The supplementary angle is 15 degrees less than 4 times the complementary angle. Find the angle.
In the office, we have come up with two different solutions: 55 degrees and 39 degrees.
Which one is right?

Answer by stanbon(75887)

The supplementary angle is 15 degrees less than 4 times the complementary angle. Find the angle.
---------------
Let the angle be "x".
The complement is "90-x"
The supplement is "180-x"
------------------
Equation:
180-x = 4(90-x)-15
180-x = 360-4x-15
3x = 165
x = 55 degrees (this is "the angle")
=====================
Cheers,
Stan H.